Consider the exponential probability density function:

for x ≥ 0

  1. Choose the correct formula for P (xx0).
    (1) P (xx0) = 1 - e(-x0/3)
    (2) P (xx0) = 1 + e(-x0/3)
    (3) P (xx0) = 1 - (1/3)e(-x0/3)
    (4) P (xx0) = 1 + (1/3)e(-x0/3)


  2. Find P (x ≤ 2). If required, round your answer to four decimal places.

    P (x ≤ 2) =
  3. Find P (x ≥ 3). If required, round your answer to four decimal places.

    P (x ≥ 3) =
  4. Find P (x ≤ 5). If required, round your answer to four decimal places.

    P (x ≤ 5) =
  5. Find P (2 ≤ x ≤ 5). If required, round your answer to four decimal places.

    P (2 ≤ x ≤ 5) =
Answer #1

(a) The exponential probability distribution is given by 1 − e−λx

This gives the area under the probability distribution function

Hence, correct option is (1) P (x ≤ x0) = 1 - e(-x0/3)

(b) P (x ≤ 2) = 1 - e(-2/3) = 0.4866

(c) P (x ≥ 3) = 1 - P (x ≤ 3) = 1 - (1 - e(-3/3)) = e-1 = 0.3679

(d) P (x ≤ 5) = 1 - e(-5/3) = 0.8111

(e) P (2 ≤ x ≤ 5) = (x ≤ 5) - (x ≤ 2) = (1 - e(-5/3) ) - (1 - e(-2/3) ) = 0.8111 - 0.4866 = 0.3245